The heat of fusion ΔHf of toluene C6H5CH3 is 6.6 /kJmol . Calculate the change in entropy

ΔS when 46.g of toluene melts at −95.0°C

Be sure your answer contains a unit symbol. Round your answer to

2 significant digits.
Answer #1

Molar mass of C7H8,

MM = 7*MM(C) + 8*MM(H)

= 7*12.01 + 8*1.008

= 92.134 g/mol

mass(C7H8)= 46 g

use:

number of mol of C7H8,

n = mass of C7H8/molar mass of C7H8

=(46 g)/(92.13 g/mol)

= 0.4993 mol

Given:

ΔH = 6.6 KJ/mol

= 6600 J/mol

use:

Q = ΔH * number of mol

= 6600 J/mol * 0.4993 mol

= 3295.2005 J

Given:

Q = 3295.2005 J

T= -95.0 oC

= (-95.0+273) K

= 178 K

use:

ΔS = Q/T

= 3295.2005 J /178 K

= 18.5124 J/K

Answer: 19 J/K